6. A particle mov es at a constant speed in a circular path with a radius of 0.02 m. If the parti acceleration? a. 20 m/s2 b. 18 m/s cle makes four revolutions each second, what is the magnitude of its c. 13 m/s d. 24 m/s2
Answer #1

number of revolutions per second = 4

radius of circle = 0.02 m

time period = 1/4 = 0.25 seconds

angular speed becomes w= (2xpi)/T = 25.12 radians/second

Angular acceleration is = rw^2 = 25.122*0.020 = 13 m/s2

Answer is [C]

Hope you will find the answer useful. Please rate. Cheers.